# Bounding the conditional von-Neumann entropy for device independent cryptography and randomness extraction

### v1

@@ V Acknowledgements
From the tools developed in [32], we use the following estimate with the definition
$y_{k}:=\sup_{0\leq P\leq\mathds{1}}\operatorname{tr}[P(\sigma t_{k}-\rho)]$ (26)
and then
@@
$\int_{\mu}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\sigma s-\rho]\leq y_{1}[(1+\frac{t_{1}}{t_{2}-t_{1}})\log\frac{t_{2}}{t_{1}}-1]+y_{r}[1-\frac{t_{r-1}}{t_{r}-t_{r-1}}\log\frac{t_{r}}{t_{r-1}}]$ (27)
$\hskip 142.26378pt+\sum_{k=2}^{r-1}y_{k}[(1+\frac{t_{k}}{t_{k+1}-t_{k}})\log\frac{t_{k+1}}{t_{k}}-\frac{t_{k-1}}{t_{k}-t_{k-1}}\log\frac{t_{k}}{t_{k-1}}].$
It is true for all $\mu>0$ and can be summarized with the following coefficients
$\alpha_{k}=\begin{cases}-[(1+\frac{t_{1}}{t_{2}-t_{1}})\log\frac{t_{2}}{t_{1}}-1]&k=1\\ -[1-\frac{t_{r-1}}{t_{r}-t_{r-1}}\log\frac{t_{r}}{t_{r-1}}]&k=r\\ -[(1+\frac{t_{k}}{t_{k+1}-t_{k}})\log\frac{t_{k+1}}{t_{k}}-\frac{t_{k-1}}{t_{k}-t_{k-1}}\log\frac{t_{k}}{t_{k-1}}]&\text{else}.\end{cases}$
and for $\beta_{k}$
$\beta_{k}=\begin{cases}[(1+\frac{t_{1}}{t_{2}-t_{1}})\log\frac{t_{2}}{t_{1}}-1]\cdot t_{1}&k=1\\ [1-\frac{t_{r-1}}{t_{r}-t_{r-1}}\log\frac{t_{r}}{t_{r-1}}]\cdot t_{r}&k=r\\ [(1+\frac{t_{k}}{t_{k+1}-t_{k}})\log\frac{t_{k+1}}{t_{k}}-\frac{t_{k-1}}{t_{k}-t_{k-1}}\log\frac{t_{k}}{t_{k-1}}]\cdot t_{k}&\text{else}.\end{cases}$
For the interval $[0,\mu]$, we use the convexity of the function $s\mapsto\operatorname{tr}^{+}[\sigma s-\rho]$ from [32, Lem. 1] and the fact that for positive operators $\sigma,\rho\in\mathcal{P}(\mathcal{H})$, the function $s\mapsto\operatorname{tr}^{+}[\sigma s-\rho]$ has the limit $\operatorname{tr}^{+}[\sigma s-\rho]\to 0$ as $s\to 0$. Thus, a simple convex interpolation on the interval $[0,\mu]$ gives in addition the following relations: $\alpha_{0}=-1$ and $\beta_{0}=\mu$. Inserting now the definition of $y_{k}$ from (26) into (27) and simplifying everything with the definitions for $\alpha_{k}$ and $\beta_{k}$ for $0\leq k\leq r$ yields the following upper estimate for the relative entropy
@@
$D(\rho\|\sigma)$ $\leq\frac{1}{\ln 2}(\operatorname{tr}[\rho-\sigma]+\sup_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\sum_{k=0}^{r}\operatorname{tr}[P_{k}(\alpha_{k}\rho+\beta_{k}\sigma)]+\ln\lambda+1-\lambda),$ (28)
which is exactly (12) and thus the assertion.
@@
For some positive operator $P_{k}^{(a)}$. Calculating $(P_{k}^{(a)})^{2}$ then yields that even the $P_{k}^{(a)}$ have to be projections if $P_{k}$ is a projection. In addition we use the following calculation for $\rho_{AQ_{E}|\tilde{x}}$
@@
Now using the fact that $\rho_{AQ_{E}|\tilde{x}}\leq\mathds{1}_{A}\otimes\rho_{Q_{E}}$ (due to the fact that the state is classical quantum) and thus $\lambda=1$ yields $\ln\lambda+1-\lambda=0$. Moreover, as discussed in Theorem II.1 and the proof in Appendix A, $\mu=0$ can be handled without problems such that the constraints coming from the integral representation in (5) become trivial. Moreover, the operators $P_{k}^{(a)}\in\mathcal{B}(\mathcal{H}_{Q_{E}})$ such that they commute with all local measurements on Alice and Bob’s side respectively. Adding the fact that they are projections from (31), we conclude all the constraints stated in Theorem II.2. Observing the purity argument of $\rho_{Q_{A}Q_{B}Q_{E}}$ by Naimark’s Theorem (e.g. [34, Chap. Dilation Theory]) yields the assertion of the Theorem II.2, if we replace tensor products with commuting operators.
@@ Appendix C Global randomness from the full distribution
$=\sum_{a\in A,b\in B}|a,b\rangle\langle a,b|\otimes\operatorname{tr}_{Q_{A}Q_{B}}(\rho_{Q_{A}Q_{B}Q_{E}}M_{a|\tilde{x}}\otimes N_{b|\tilde{y}}).$
@@ Appendix D Algebraic Perspective on the NPA hierarchy
An illustrative example of this methodology is the sum-of-squares (SOS) cones, which were initially introduced in [35, 37]. For SOS cones, the Python package described in [39] automates the translation of generators and relations into an SDP within a matrix algebra framework.
@@ Appendix E Additional discussion for the numerics
along with some associated constraints. As discussed in [3, Rem. 2.6], we can implement several optimizations to accelerate the numerical evaluations:

### v2

@@ VIII Acknowledgements
The following is a generalization of [25, Cor. 1] to positive trace-class operators.
Lemma A.1.
Let $\rho,\sigma\in\mathcal{P}(\mathcal{H})$ be positive trace-class operators such that there exist $\mu,\lambda\geq 0$ such that $\mu\,\sigma\leq\rho\leq\lambda\,\sigma$. Then
$D(\rho\|\sigma)=\operatorname{tr}[\rho-\sigma]+\int_{\mu}^{\lambda}\frac{ds}{s}\,\operatorname{tr}^{-}\!\big{[}\rho-s\sigma\big{]}+\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma].$ (31)
Proof.
We start from the integral representation of relative entropy valid for all positive trace-class operators (see [18]):
$D(\rho\|\sigma)=\operatorname{tr}[\rho-\sigma]+\int_{-\infty}^{\infty}\frac{|t|}{(1-t)^{2}}\,\operatorname{tr}^{-}\!\big{[}(1-t)\rho+t\sigma\big{]}\,dt.$ (32)
Split the integral at $t=0$ and $t=1$. For $t\leq 0$ we have 1t>01-t>0 and
$((1-t)\rho+t\sigma)_{-}=(1-t)\bigl{(}\rho-\tfrac{t}{t-1}\sigma\bigr{)}_{-}.$ (33)
With the substitution $s=\tfrac{t}{t-1}$ (which maps $(-\infty,0]\to[0,1]$), one checks that
$\int_{-\infty}^{0}\frac{|t|}{(1-t)^{2}}\operatorname{tr}^{-}[(1-t)\rho+t\sigma]\,dt=\int_{0}^{1}\frac{ds}{s}\,\operatorname{tr}^{-}[\rho-s\sigma].$ (34)
Since $\mu\sigma\leq\rho$, the integrand vanishes for $s\in[0,\mu]$, hence the last integral equals $\int_{\mu}^{1}\frac{ds}{s}\operatorname{tr}^{-}[\rho-s\sigma]$.
For $t\geq 1$, we use
$((1-t)\rho+t\sigma)_{-}=((t-1)\rho-t\sigma)_{+}=(t-1)\bigl{(}\rho-\tfrac{t}{t-1}\sigma\bigr{)}_{+}.$ (35)
With the same substitution $s=\tfrac{t}{t-1}$ (mapping $1,\infty)\to[1,\infty)$),
$\int_{1}^{\infty}\frac{t}{(1-t)^{2}}\operatorname{tr}^{-}[(1-t)\rho+t\sigma]\,dt=\int_{1}^{\infty}\frac{ds}{s}\,\operatorname{tr}^{+}[\rho-s\sigma].$ (36)
@@
Because $\rho\leq\lambda\sigma$, we have $(\rho-s\sigma)_{+}=0$ for all $s\geq\lambda$, so the integral truncates to $\int_{1}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\rho-s\sigma]$.
Insert these two pieces in (32) to get
$D(\rho\|\sigma)=\operatorname{tr}[\rho-\sigma]+\int_{\mu}^{1}\frac{ds}{s}\operatorname{tr}^{-}[\rho-s\sigma]+\int_{1}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\rho-s\sigma].$ (37)
Now use the elementary identity for any self-adjoint $X$: $\operatorname{tr}[X_{+}]=\operatorname{tr}[X_{-}]+\operatorname{tr}[X].$ With $X=\rho-s\sigma$ this gives
$\operatorname{tr}^{+}[\rho-s\sigma]=\operatorname{tr}^{-}[\rho-s\sigma]+\operatorname{tr}[\rho-s\sigma].$ (38)
Hence
$\int_{1}^{\lambda}\frac{ds}{s}\operatorname{tr}^{+}[\rho-s\sigma]=\int_{1}^{\lambda}\frac{ds}{s}\operatorname{tr}^{-}[\rho-s\sigma]+\int_{1}^{\lambda}\frac{ds}{s}\bigl{(}\operatorname{tr}[\rho]-s\,\operatorname{tr}[\sigma]\bigr{)}.$ (39)
Combining with the $\mu,1]$ piece collapses the two $\operatorname{tr}^{-}[\rho-s\sigma]$-integrals into a single integral over $\mu,\lambda]$, and the elementary integral evaluates to
$\int_{1}^{\lambda}\frac{ds}{s}\bigl{(}\operatorname{tr}[\rho]-s\,\operatorname{tr}[\sigma]\bigr{)}=\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma].$ (40)
This yields (31).
Finally, independence of the choice of $\mu,\lambda$ follows from the truncation properties: for $s\leq\mu$ we have $(\rho-s\sigma)_{-}=0$, and for $s\geq\lambda$ we have $(\rho-s\sigma)_{+}=0$. Equivalently, if $\lambda^{\prime}\geq\lambda$ then on $(\lambda,\lambda^{\prime}]$ one has $(\rho-s\sigma)_{-}=s\,\operatorname{tr}[\sigma]-\operatorname{tr}[\rho]$, so the change in $\int\frac{ds}{s}\operatorname{tr}^{-}[\rho-s\sigma]$ is exactly canceled by the change in $\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma]$, and an analogous statement holds at the lower limit $\mu$. ∎
Proof of Theorem II.1:
Proof.
Fix a partition $0<t_{1}<\cdots<t_{r}$ with $t_{1}=\mu$ and $t_{r}=\lambda$. Following the tools in [28], set
$y_{k}:=\sup_{0\leq P\leq\mathds{1}}\operatorname{tr}\big{[}P(\sigma t_{k}-\rho)\big{]}.$ (41)
Then [28] yields the estimate
$\int_{\mu}^{\lambda}\frac{ds}{s}\,\operatorname{tr}^{+}[\sigma s-\rho]$ $\leq y_{1}\Big{[}\Big{(}1+\frac{t_{1}}{t_{2}-t_{1}}\Big{)}\ln\frac{t_{2}}{t_{1}}-1\Big{]}+y_{r}\Big{[}1-\frac{t_{r-1}}{t_{r}-t_{r-1}}\ln\frac{t_{r}}{t_{r-1}}\Big{]}$ (42)
$\qquad+\sum_{k=2}^{r-1}y_{k}\Big{[}\Big{(}1+\frac{t_{k}}{t_{k+1}-t_{k}}\Big{)}\ln\frac{t_{k+1}}{t_{k}}-\frac{t_{k-1}}{t_{k}-t_{k-1}}\ln\frac{t_{k}}{t_{k-1}}\Big{]}.$
It is convenient to encode the coefficients by
$\alpha_{k}:=\begin{cases}-\Big{[}\Big{(}1+\dfrac{t_{1}}{t_{2}-t_{1}}\Big{)}\ln\dfrac{t_{2}}{t_{1}}-1\Big{]}&k=1,\\[4.0pt] -\Big{[}1-\dfrac{t_{r-1}}{t_{r}-t_{r-1}}\ln\dfrac{t_{r}}{t_{r-1}}\Big{]}&k=r,\\[6.0pt] -\Big{[}\Big{(}1+\dfrac{t_{k}}{t_{k+1}-t_{k}}\Big{)}\ln\dfrac{t_{k+1}}{t_{k}}-\dfrac{t_{k-1}}{t_{k}-t_{k-1}}\ln\dfrac{t_{k}}{t_{k-1}}\Big{]}&\text{else},\end{cases}$
and
$\beta_{k}:=\begin{cases}\Big{[}\Big{(}1+\dfrac{t_{1}}{t_{2}-t_{1}}\Big{)}\ln\dfrac{t_{2}}{t_{1}}-1\Big{]}\,t_{1}&k=1,\\[4.0pt] \Big{[}1-\dfrac{t_{r-1}}{t_{r}-t_{r-1}}\ln\dfrac{t_{r}}{t_{r-1}}\Big{]}\,t_{r}&k=r,\\[6.0pt] \Big{[}\Big{(}1+\dfrac{t_{k}}{t_{k+1}-t_{k}}\Big{)}\ln\dfrac{t_{k+1}}{t_{k}}-\dfrac{t_{k-1}}{t_{k}-t_{k-1}}\ln\dfrac{t_{k}}{t_{k-1}}\Big{]}\,t_{k}&\text{else}.\end{cases}$
For the interval $0,\mu]$, use the convexity of $s\mapsto\operatorname{tr}^{+}[\sigma s-\rho]$ [28, Lem. 1] and the fact that $\operatorname{tr}^{+}[\sigma s-\rho]\to 0$ as $s\to 0$ for positive $\sigma,\rho\in\mathcal{P}(\mathcal{H})$. For $s\in[0,\mu]$,
$\operatorname{tr}^{+}[\sigma s-\rho]\leq\Big{(}1-\frac{s}{\mu}\Big{)}\operatorname{tr}^{+}[\sigma\cdot 0-\rho]+\frac{s}{\mu}\operatorname{tr}^{+}[\sigma\mu-\rho]=\frac{s}{\mu}\operatorname{tr}^{+}[\sigma\mu-\rho].$
Hence
$\int_{0}^{\mu}\frac{ds}{s}\,\operatorname{tr}^{+}[\sigma s-\rho]$ $\leq\operatorname{tr}^{+}[\sigma\mu-\rho]=\sup_{0\leq P\leq\mathds{1}}\operatorname{tr}\big{[}P(\sigma\mu-\rho)\big{]},$
which corresponds to the choice $\alpha_{0}=-1$ and $\beta_{0}=\mu$.
Inserting (41) into (42), and using the definitions of $\alpha_{k},\beta_{k}$ for $0\leq k\leq r$, we obtain with A.1
@@
$D(\rho\|\sigma)\leq\frac{1}{\ln 2}\Big{(}\operatorname{tr}[\rho-\sigma]+\sup_{0\leq P_{0},\ldots,P_{r}\leq\mathds{1}}\sum_{k=0}^{r}\operatorname{tr}\big{[}P_{k}(\alpha_{k}\rho+\beta_{k}\sigma)\big{]}+\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma]\Big{)},$
which is exactly (13). This proves the claim. ∎
@@
For some positive operator $P_{k}^{(a)}$. Calculating $(P_{k})^{2}$ then yields that even the $P_{k}^{(a)}$ have to be projections if $P_{k}$ is a projection. In addition we use the following calculation for $\rho_{AQ_{E}|\tilde{x}}$
@@
Now using the fact that $\rho_{AQ_{E}|\tilde{x}}\leq\mathds{1}_{A}\otimes\rho_{Q_{E}}$ (due to the fact that the state is classical quantum) and thus $\lambda=1$ yields $\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma]=0$. Moreover, as discussed in Theorem II.1 and the proof in Appendix A, $\mu=0$ can be handled without problems such that the constraints coming from the integral representation in (6) become trivial. Moreover, the operators $P_{k}^{(a)}\in\mathcal{B}(\mathcal{H}_{Q_{E}})$ such that they commute with all local measurements on Alice and Bob’s side respectively. Adding the fact that they are projections from (45), we conclude all the constraints stated in Theorem II.2. Observing the purity argument of $\rho_{Q_{A}Q_{B}Q_{E}}$ by Naimark’s Theorem yields the assertion of the Theorem II.2, if we replace tensor products with commuting operators. ∎
@@ Appendix C Global randomness from the full distribution
[…]
@@ Appendix D Algebraic Perspective on the NPA hierarchy
An illustrative example of this methodology is the sum-of-squares (SOS) cones, which were initially introduced in [19]. For SOS cones, the Python package described in [31] automates the translation of generators and relations into an SDP within a matrix algebra framework.
@@ Appendix E Additional discussion for the numerics
along with some associated constraints. As discussed in [22, Rem. 2.6], we can implement several optimizations to accelerate the numerical evaluations: